# Lower bounding the MaxCut of high girth 3-regular graphs using the QAOA

### v1

@@ 1 Introduction
We use the infimum because the set of graphs $G$ with girth($G$)$\geq g$ is infinite, but (1) means that every graph $G$ of girth at least $g$ has a cut of size at least $|E(G)|M_{g}$. The numerical value of $M_{g}$ is currently unknown for general $g$.
@@ 3 Pre-computing QAOA parameters for graphs with large girth
$F_{p}({\boldsymbol{\gamma}},{\boldsymbol{\beta}})=|E(G)|\leavevmode\nobreak\ c_{\rm edge}({\boldsymbol{\gamma}},{\boldsymbol{\beta}})$ (7)
@@
$\tilde{c}_{\rm edge}(p)=c_{\rm edge}(\boldsymbol{\tilde{\gamma}},\boldsymbol{\tilde{\beta}}).$ (10)
@@
When running the QAOA for graphs of girth greater than 36, it would be helpful to pre-compute the optimal parameters for $p>17$, which we have not done. For sufficiently large $p$ this may be intractable, or it may be well-approximated by extrapolation of the optimal parameters from smaller $p$ to larger $p$. Regardless, for the purpose of proposing a formal algorithm, for all 3-regular graphs of girth greater than 36, we propose running the $p=17$ QAOA with our fixed, pre-computed parameters. This protocol already gives an algorithm with polynomial runtime with respect to graph size, with the best currently known performance guarantee on cut fraction. The only other algorithms we currently know with the same performance guarantee take exponential time. One example is the the brute force algorithm, which finds the best cut in exponential time, and is therefore guaranteed to find a cut of size at least $0.8971|E(G)|$, whose existence follows from our present results. A classical algorithm of Williams [13] finds the best cut in exponential time but using a smaller exponent than brute force search.
@@ 4 Analyzing time-complexity
Figure 3: Optimized parameters $\boldsymbol{\tilde{\gamma}}$ and $\boldsymbol{\tilde{\beta}}$ as a function of $(j-1)/(p-1)$ for different values of $p$ up to p=17. These figures suggest that the optimal parameters might approach fixed curves as $p$ increases.
@@
After running the QAOA circuit and performing a measurement in the computational basis, we obtain a cut of the graph that depends on the measurement outcome. The expected value of the cut size is $|E(G)|\tilde{c}_{\rm edge}(p)$. We can also ask for a guarantee that the cut is above (the floor of) this expected value with high probability. To that end, repeat the QAOA circuit and measurement many times and record the best outcome. The probability of sampling a cut size of at least $\lfloor|E(G)|\tilde{c}_{\rm edge}(p)\rfloor$ with a single sample is at least $1/|E(G)|$. So with $O(|E(G)|)$ repetitions, we obtain a sample with cut at least $\lfloor|E(G)|\tilde{c}_{\rm edge}(p)\rfloor$, with probability at least $2/3$ (or any constant probability below 1).
@@ 5.1 Results
In order to maximize the quantum expected value of the cut size, we numerically evaluate the quantum expectation value of the cost function of Eq. (4). As we emphasized before, at the QAOA depths we consider, on these high girth graphs, each edge has an isomorphic tree neighborhood and so each edge gives the same contribution. We find values of $({\boldsymbol{\gamma}},{\boldsymbol{\beta}})$ that approximately maximize $c_{\rm edge}({\boldsymbol{\gamma}},{\boldsymbol{\beta}})$. For fixed $({\boldsymbol{\gamma}},{\boldsymbol{\beta}})$ the quantum expectation is evaluated using the tensor network contraction method described in Section 5.2. The blue crosses in Fig. 2 show the optimized values of $c_{\rm edge}({\boldsymbol{\gamma}},{\boldsymbol{\beta}})$ at approximately optimal parameters, $\tilde{c}_{\rm edge}(p)=c_{\rm edge}(\boldsymbol{\tilde{\gamma}},\boldsymbol{\tilde{\beta}})$. The values are listed in Table 1. The parameters $\boldsymbol{\tilde{\gamma}}$ and $\boldsymbol{\tilde{\beta}}$ are shown in Fig. 3 for different values of $p$.
@@ 5.2 Methods
[Uncaptioned image] $=\quad\frac{1}{\sqrt{2}}\quad\forall a$ [Uncaptioned image] =$=\quad\begin{cases}e^{i\gamma}&{\rm if}\,\,a=b\\ e^{-i\gamma}&{\rm if}\,\,a\neq b\end{cases}$
[Uncaptioned image] =$=\quad\begin{cases}\cos{(\beta)}&{\rm if}\,\,a=b\\ -i\sin{(\beta)}&{\rm if}\,\,a\neq b\end{cases}$ [Uncaptioned image] =$=\quad\begin{cases}1&{\rm if}\,\,a=b=0\\ -1&{\rm if}\,\,a=b=1\\ 0&{\rm otherwise}\end{cases}$ (12)
@@
Figure 4: We plot $\tilde{c}_{\rm edge}$ which is $c_{\rm edge}$ at optimized angles $\boldsymbol{\tilde{\gamma}}$ and $\boldsymbol{\tilde{\beta}}$ as a function of $1/p$. The value 0.912 is the lower bound on $\lim_{g\to\infty}M_{g}$ given by Refs. [6, 5]. If $\tilde{c}_{\rm edge}$ were to exceed this value at larger $p$, it would provide a new lower bound on $\lim_{g\to\infty}M_{g}$. The value 0.9351 is the upper bound on the expected cut fraction of large random $3$-regular graphs given in Refs. [16, 17], and $\tilde{c}_{\rm edge}$ cannot exceed this value.
In the generic case of $d>2$ the number of qubits in the light-cone of the $Z_{i}Z_{j}$ operator grows exponentially with $p$. For a 3-regular graph (i.e. $d=3$), at $p=1,2,3,\ldots$ the size of the light cone is $6,14,30,\ldots$ In general, for a $d$-regular graph, the number of qubits at depth $p$ is $2\frac{(d-1)^{p+1}-1}{d-2}$. For $d=3$ and $p=17$ the size of the light-cone is 524,286 qubits. However, due to the fact that all branches in the regular tree are identical, we can perform this computation in a compact way that is not affected by the exponential growth in the number of qubits in the calculation. We contract a single branch inwards towards the root of the tree, raising its tensor entries to the $(d-1)$th power before proceeding to the next level of the tree. This is expressed in graphical notation in Fig. 5(b), where we have made use of the definition
[Uncaptioned image]$\equiv\quad\left(T_{a_{1},a_{2},\ldots,a_{n}}\right)^{d-1}$ (13)
@@
Figure 5: Tensor network for the computation of $\langle{\boldsymbol{\gamma}},{\boldsymbol{\beta}}|Z_{i}Z_{j}|{\boldsymbol{\gamma}},{\boldsymbol{\beta}}\rangle$ for $p=3$. On the left (a) we have $d=2$ and on the right (b) arbitrary $d$. In the latter case, the tensor network diagramatic notation is extended to denote extry-wise exponentiation of a tensor. In particular, the result of contracting the tensors in one of the colored boxes is raised to the power $(d-1)$ before proceeding to later contractions, as expressed in Eq. (13). The time and space complexities of the contraction performed in this way are both $\mathcal{O}(2^{2p})$. The complexity does not depend on $d$. Higher values of $p$ are tackled in a similar fashion.
@@ 6 MaxCut Comparisons
The graphs which we have considered have large girth and every edge sits in a tree neighborhood. We can ask about large random 3-regular graphs where almost all edges sit in tree neighborhoods. Such graphs have an order unity number of triangles, squares, pentagons, or any cycle of constant length, which means that they are not strictly large girth. However, since there are so few short cycles, any lower bound on $M_{g}$ for any $g$ is a lower bound on the cut fraction of a typical large random 3-regular graph [6]. Furthermore the QAOA at level $p$ achieves cut fraction $\tilde{c}_{\rm edge}(p)$ on these graphs. There is an upper bound on the cut fraction of typical large random 3-regular graphs of 0.9351, described in Ref. [21] and shown by Hladky and McKay [16, 17]. Therefore the QAOA cut fraction $\tilde{c}_{\rm edge}(p)$ cannot exceed 0.9351 even as $p\to\infty$. This value is shown in Fig. 4.
@@ 7 Maximum Independent Set
written as a function of the bit string $(b_{1},\ldots,b_{|V|})$, where each bit string also represents a subset of the vertex set $V$, with $b_{i}=1$ if vertex $i$ is in the subset. Note this cost function is defined on all subsets and not just independent sets. The first term of $I_{1}$ is the Hamming weight which we want to make big, and the subtracted term counts violations of the independent set property. We now argue that given any bit string with associated cost $I_{1}$, there exists an independent set of size at least $I_{1}$. In particular, we consider the set of vertices labeled 1 in the bit string, and we construct a subset of these vertices that forms an independent set of size at least $I_{1}$. To begin, given a bit string with cost $I_{1}$, assume that the set of $1$’s is not yet an independent set. Choose any violated edge $\langle i,j\rangle$, that is with $b_{i}=b_{j}=1$, then choose either $i$ or $j$ and re-label it as 00, that is, remove the vertex from the set under consideration. Consider the cost function $I_{1}$ for this new string: the bit flip decreases the Hamming weight by $1$ and decreases the number of violations by at least $1$, so $I_{1}$ cannot decrease. Repeat this until there are no more violations so the final bit string must have $1$’s that form a valid independent set. The new value of $I_{1}$ is equal the size of the independent set which must be at least the original value of $I_{1}$. Now the maximum of $I_{1}$ occurs when $b$ corresponds to an actual maximum independent set so achieving a high value of $I_{1}$ gives an approximation to the size of the largest independent set.
@@
Write the two terms as $I_{2}=I_{2,1}+I_{2,0}$. For any input bit string, the $1$’s represent one set and the 00’s the complementary set. (Note that 00 and $1$ here do not mean what they meant in the previous paragraph.) By the above argument regarding $I_{1}$, first applied to $I_{2,1}$, there exists a subset of $1$’s that forms an independent set, with size at least $I_{2,1}$. Meanwhile, the same argument applied to $I_{2,0}$ implies there exists a subset of 00’s that forms an independent set, with size at least $I_{2,0}$. Thus we have two disjoint independent sets of total size at least $I_{2}$.
@@
Figure 7: Optimized parameters $\boldsymbol{\tilde{\gamma}}$, $\boldsymbol{\tilde{\gamma^{\prime}}}$ and $\boldsymbol{\tilde{\beta}}$ as a function of $(j-1)/(p-1)$ for different values of $p$ up to p=15 for the MIS problem.
In Fig. 6 we show the lower bounds we get on the independence ratio as a function of $1/p$ from our MaxCut results and using the $3$-parameter per layer ansatz. The $3$-parameter ansatz must lie above the other but the difference shrinks as $p$ grows. The red line is the best available bound on the independence ratio for large girth 3-regular graphs as the girth goes to infinity [22]. The optimized parameters $(\tilde{{\boldsymbol{\gamma}}},\tilde{\boldsymbol{\gamma^{\prime}}},\tilde{{\boldsymbol{\beta}}})$ are shown in Fig. 7 for different values of $p$.

### v2

@@ 1 Introduction
[…]
@@ 3 Pre-computing QAOA parameters for graphs with large girth
$F_{p}({\boldsymbol{\gamma}},{\boldsymbol{\beta}})=|E(G)|\penalty 10000\ c_{\rm edge}({\boldsymbol{\gamma}},{\boldsymbol{\beta}})$ (7)
@@
[…]
@@
When running the QAOA for graphs of girth greater than 36, it would be helpful to pre-compute the optimal parameters for $p>17$, which we have not done. For sufficiently large $p$ this may be intractable, or it may be well-approximated by extrapolation of the optimal parameters from smaller $p$ to larger $p$. Regardless, for the purpose of proposing a formal algorithm, for all 3-regular graphs of girth greater than 36, we propose running the $p=17$ QAOA with our fixed, pre-computed parameters. This protocol already gives an algorithm with polynomial runtime with respect to graph size, with the best currently known performance guarantee on cut fraction. The only other algorithms we currently know with the same performance guarantee take exponential time. One example is the brute force algorithm, which finds the best cut in exponential time, and is therefore guaranteed to find a cut of size at least $0.8971|E(G)|$, whose existence follows from our present results. A classical algorithm of Williams [20] finds the best cut in exponential time but using a smaller exponent than brute force search.
@@ 4 Analyzing time-complexity
Figure 3: Optimized parameters $\boldsymbol{\tilde{\gamma}}$ and $\boldsymbol{\tilde{\beta}}$ as a function of (j1)/(p1)(j-1)/(p-1) for different values of $p$ up to p=17. These figures suggest that the optimal parameters might approach fixed curves as $p$ increases.
@@
After running the QAOA circuit and performing a measurement in the computational basis, we obtain a cut of the graph that depends on the measurement outcome. The expected value of the cut size is $|E(G)|\tilde{c}_{\rm edge}(p)$. We can also ask for a guarantee that the cut is above (the floor of) this expected value with high probability. To that end, repeat the QAOA circuit and measurement many times and record the best outcome. The probability of sampling a cut size of at least $\lfloor|E(G)|\tilde{c}_{\rm edge}(p)\rfloor$ with a single sample is at least $1/|E(G)|$. To see this, let $C\in\{0,1,\ldots,|E(G)|\}$ denote the cut size obtained from a single sample, and denote $C_{0}=\lfloor|E(G)|\,\tilde{c}_{\rm{edge}}(p)\rfloor$, with $p_{0}:=\Pr\!\big(C\geq C_{0}\big)$. We calculate
$\begin{split}\mathbb{E}[C]&\leq(C_{0}-1)\Pr(C\leq C_{0}-1)+|E(G)|\,\Pr(C\geq C_{0})\\ &=(1-p_{0})(C_{0}-1)+p_{0}\,|E(G)|\\ &\leq C_{0}-1+p_{0}|E(G)|,\end{split}$ (12)
where the first line uses the fact $C$ takes integer values with $C\leq|E(G)|$. Combining the above with $\mathbb{E}[C]\geq C_{0}$, we obtain $p_{0}\geq 1/|E(G)|$. Then with $O(|E(G)|)$ repetitions of the QAOA circuit and measurement, we obtain a sample with cut at least $\lfloor|E(G)|\tilde{c}_{\rm edge}(p)\rfloor$, with probability at least $2/3$ (or any constant probability below 1).
@@ 5.1 Results
[…]
@@ 5.2 Methods
[Uncaptioned image] $=\quad\frac{1}{\sqrt{2}}\quad\forall a$ [Uncaptioned image] $=\quad\begin{cases}e^{i\gamma}&{\rm if}\,\,a=b\\ e^{-i\gamma}&{\rm if}\,\,a\neq b\end{cases}$
[Uncaptioned image] $=\quad\begin{cases}\cos{(\beta)}&{\rm if}\,\,a=b\\ -i\sin{(\beta)}&{\rm if}\,\,a\neq b\end{cases}$ [Uncaptioned image] $=\quad\begin{cases}1&{\rm if}\,\,a=b=0\\ -1&{\rm if}\,\,a=b=1\\ 0&{\rm otherwise}\end{cases}$ (13)
@@
Figure 4: We plot $\tilde{c}_{\rm edge}$ which is $c_{\rm edge}$ at optimized angles $\boldsymbol{\tilde{\gamma}}$ and $\boldsymbol{\tilde{\beta}}$ as a function of $1/p$. The value 0.912 is the lower bound on $\lim_{g\to\infty}M_{g}$ given by Refs. [4, 8]. If $\tilde{c}_{\rm edge}$ were to exceed this value at larger $p$, it would provide a new lower bound on $\lim_{g\to\infty}M_{g}$. The value 0.9239 is the upper bound on the expected cut fraction of large random 33-regular graphs given in [12], and $\tilde{c}_{\rm edge}$ cannot exceed this value.
In the generic case of $d>2$ the number of qubits in the light-cone of the $Z_{i}Z_{j}$ operator grows exponentially with $p$. For a 3-regular graph (i.e. $d=3$), at $p=1,2,3,\ldots$ the size of the light cone is $6,14,30,\ldots$ In general, for a $d$-regular graph, the number of qubits at depth $p$ is $2\frac{(d-1)^{p+1}-1}{d-2}$. For $d=3$ and $p=17$ the size of the light-cone is 524,286 qubits. However, due to the fact that all branches in the regular tree are identical, we can perform this computation in a compact way that is not affected by the exponential growth in the number of qubits in the calculation. We contract a single branch inwards towards the root of the tree, raising its tensor entries to the (d1)(d-1)th power before proceeding to the next level of the tree. This is expressed in graphical notation in Fig. 5(b), where we have made use of the definition
[…]
@@
Figure 5: Tensor network for the computation of $\langle{\boldsymbol{\gamma}},{\boldsymbol{\beta}}|Z_{i}Z_{j}|{\boldsymbol{\gamma}},{\boldsymbol{\beta}}\rangle$ for $p=3$. On the left (a) we have $d=2$ and on the right (b) arbitrary $d$. In the latter case, the tensor network diagramatic notation is extended to denote extry-wise exponentiation of a tensor. In particular, the result of contracting the tensors in one of the colored boxes is raised to the power (d1)(d-1) before proceeding to later contractions, as expressed in Eq. (14). The time and space complexities of the contraction performed in this way are both $\mathcal{O}(2^{2p})$. The complexity does not depend on $d$. Higher values of $p$ are tackled in a similar fashion.
@@ 6 MaxCut Comparisons
The graphs which we have considered have large girth and every edge sits in a tree neighborhood. We can ask about large random 3-regular graphs where almost all edges sit in tree neighborhoods. Such graphs have an order unity number of triangles, squares, pentagons, or any cycle of constant length, which means that they are not strictly large girth. However, since there are so few short cycles, any lower bound on $M_{g}$ for any $g$ is a lower bound on the cut fraction of a typical large random 3-regular graph [8]. Furthermore the QAOA at level $p$ achieves cut fraction $\tilde{c}_{\rm edge}(p)$ on these graphs. There is an upper bound on the cut fraction of typical large random 3-regular graphs of 0.9239 [12], building on Refs. [3, 24, 13, 16]. Therefore the QAOA cut fraction $\tilde{c}_{\rm edge}(p)$ cannot exceed 0.9239 even as $p\to\infty$. This value is shown in Fig. 4.
@@ 7 Maximum Independent Set
written as a function of the bit string $(b_{1},\ldots,b_{|V|})$, where each bit string also represents a subset of the vertex set $V$, with $b_{i}=1$ if vertex $i$ is in the subset. Note this cost function is defined on all subsets and not just independent sets. The first term of $I_{1}$ is the Hamming weight which we want to make big, and the subtracted term counts violations of the independent set property. We now argue that given any bit string with associated cost $I_{1}$, there exists an independent set of size at least $I_{1}$. In particular, we consider the set of vertices labeled 1 in the bit string, and we construct a subset of these vertices that forms an independent set of size at least $I_{1}$. To begin, given a bit string with cost $I_{1}$, assume that the set of $1$’s is not yet an independent set. Choose any violated edge $\langle i,j\rangle$, that is with $b_{i}=b_{j}=1$, then choose either $i$ or $j$ and re-label it as 0, that is, remove the vertex from the set under consideration. Consider the cost function $I_{1}$ for this new string: the bit flip decreases the Hamming weight by $1$ and decreases the number of violations by at least $1$, so $I_{1}$ cannot decrease. Repeat this until there are no more violations so the final bit string must have $1$’s that form a valid independent set. The new value of $I_{1}$ is equal the size of the independent set which must be at least the original value of $I_{1}$. Now the maximum of $I_{1}$ occurs when $b$ corresponds to an actual maximum independent set so achieving a high value of $I_{1}$ gives an approximation to the size of the largest independent set.
@@
Write the two terms as $I_{2}=I_{2,1}+I_{2,0}$. For any input bit string, the $1$’s represent one set and the 0’s the complementary set. (Note that 0 and $1$ here do not mean what they meant in the previous paragraph.) By the above argument regarding $I_{1}$, first applied to $I_{2,1}$, there exists a subset of $1$’s that forms an independent set, with size at least $I_{2,1}$. Meanwhile, the same argument applied to $I_{2,0}$ implies there exists a subset of 0’s that forms an independent set, with size at least $I_{2,0}$. Thus we have two disjoint independent sets of total size at least $I_{2}$.
@@
Figure 7: Optimized parameters $\boldsymbol{\tilde{\gamma}}$, $\boldsymbol{\tilde{\gamma^{\prime}}}$ and $\boldsymbol{\tilde{\beta}}$ as a function of (j1)/(p1)(j-1)/(p-1) for different values of $p$ up to p=15 for the MIS problem.
[…]